OpenStudy (anonymous):

The equation of a line is -6x-2y=-18. What is the x-intercept of the line?

OpenStudy (anonymous):

-3 9 -9 3

Vocaloid (vocaloid):

when y = 0, what does x = ?

OpenStudy (anonymous):

0?

Vocaloid (vocaloid):

look at your equation

Vocaloid (vocaloid):

set y = 0 and calculate x

Vocaloid (vocaloid):

-6x - 2(0) = -18 x = __?

OpenStudy (anonymous):

I dont know..

Vocaloid (vocaloid):

-2*0 = 0 -6x = -18 x = ?

OpenStudy (anonymous):

-9?

Vocaloid (vocaloid):

check your math again... divide both sides by -6

OpenStudy (anonymous):

It not going to be negative right?

Vocaloid (vocaloid):

nope, it'll be positive

Vocaloid (vocaloid):

(-18)/(-6)=?

OpenStudy (anonymous):

3?

Vocaloid (vocaloid):

good, that's our answer

OpenStudy (anonymous):

Wait, I was actually right? wow.

OpenStudy (anonymous):

Thank you